### v2

# GI-NAS: Boosting Gradient Inversion Attacks through Adaptive Neural Architecture Search

Abstract
Gradient Inversion Attacks invert the transmitted gradients in Federated Learning (FL) systems to reconstruct the sensitive data of local clients and have raised considerable privacy concerns. A majority of gradient inversion methods rely heavily on explicit prior knowledge (e.g., a well pre-trained generative model), which is often unavailable in realistic scenarios. To alleviate this issue, researchers have proposed to leverage the implicit prior knowledge of an over-parameterized network. However, they only utilize a fixed neural architecture for all the attack settings. This would hinder the adaptive use of implicit architectural priors and consequently limit the generalizability. In this paper, we further exploit such implicit prior knowledge by proposing Gradient Inversion via Neural Architecture Search (GI-NAS), which adaptively searches the network and captures the implicit priors behind neural architectures. Extensive experiments verify that our proposed GI-NAS can achieve superior attack performance compared to state-of-the-art gradient inversion methods, even under more practical settings with high-resolution images, large-sized batches, and advanced defense strategies.
Index Terms:
Federated Learning, Gradient Inversion Attacks, Neural Architecture Search, Privacy Leakage.
I Introduction
Federated Learning (FL) [1, 2, 3, 4] serves as an efficient collaborative learning framework where multiple participants cooperatively train a global model and only the computed gradients are exchanged. By adopting this distributed paradigm, FL systems fully leverage the huge amounts of data partitioned across various clients for enhanced model efficacy and tackle the separateness of data silos [5]. Moreover, since merely the gradients instead of the intimate data are uploaded to the server, the user privacy seems to be safely guaranteed as the private data is only available at the client side.
However, FL systems are actually not so reliable as what people have expected. Extensive studies have discovered that even the conveyed gradients can disclose the sensitive information of users. Early works [6, 7, 8, 9] involve inferring the existence of certain samples in the dataset (i.e., Membership Inference Attacks [10]) or further revealing some properties of the private training set (i.e., Property Inference Attacks [11]) from the uploaded gradients. But unlike the above inference attacks that only partially reveal limited information of the private data, Gradient Inversion Attacks stand out as a more threatening privacy risk as they can completely reconstruct the sensitive data by inverting the gradients.
Zhu et al. [12] first formulate gradient inversion as an optimization problem and recover the intimate training images by minimizing the gradient matching loss (i.e., the distance between the dummy gradients and the real gradients) with image pixels regarded as trainable parameters. Ensuing works improve the attack performance on the basis of Zhu et al. [12] by designing label extraction techniques [13], switching the distance metric and introducing regularization [14], or considering larger batch sizes settings [15], but still restrict their optimizations in the pixel space. To fill this gap, recent studies [16, 17, 18] propose to explore various search algorithms within the Generative Adversarial Networks (GAN) [19, 20] to leverage the rich prior knowledge encoded in the pre-trained generative models. While incorporating these explicit priors indeed improves the attack performance, it is usually tough to pre-prepare such prerequisites in realistic FL scenarios where the data distribution at the client side is likely to be complex or unknown to the attackers. Therefore, Zhang et al. [21] propose to employ an over-parameterized convolutional network for gradient inversion and directly optimize the excessive network parameters on a fixed input, which does not require any explicit prior information but still outperforms GAN-based attacks. The reason behind this is that the structure of a convolutional network naturally captures image statics prior to any learning [22] and Zhang et al. [21] leverage this characteristic as implicit prior knowledge. However, Zhang et al. [21] only utilize a fixed network for all the attack settings, regardless of the specific batch to recover. An intuitive question naturally arises: Can we adaptively search the architecture of the over-parameterized network to better capture the implicit prior knowledge for each reconstructed batch?
(a) Quantitative Results
Original
Model $1$
Model $2$
Model $3$
Model $4$
Model $5$
(b) Qualitative Results
Figure 1: Quantitative and qualitative results when randomly attacking $3$ different batches by $5$ different models on ImageNet. In Figure 1(b), the first, second, and third rows are respectively from Batch $1$, Batch $2$, and Batch $3$.
In Figure 1, we randomly select $5$ over-parameterized networks with different architectures to attack $3$ different batches. All the networks are optimized for the same number of iterations on ImageNet [23]. The results indicate that the PSNR performance varies significantly when changing the architectures. For the same batch, various architectures can hold remarkably different implicit priors on it. Besides, since the optimal models for Batch $1$, Batch $2$, and Batch $3$ are respectively Model $1$, Model $3$, and Model $4$, there exists no network that can consistently perform the best on all the given batches. Thus, the changeless adoption of architecture by Zhang et al. [21] lacks optimality under dynamic FL scenarios where different batches require different implicit architectural priors, and it is of great significance to adaptively select the most suitable architecture for each batch.
Inspired by the above phenomenons, we propose a novel gradient inversion method, named Gradient Inversion via Neural Architecture Search (GI-NAS), to better match each batch with the optimal model architecture. Specifically, we first enlarge the potential search space for the over-parameterized network by designing different upsampling modules and skip connection patterns. To reduce the computational overhead, we utilize a training-free search strategy that compares the initial gradient matching loss for a given batch over all the candidates and selects the best of them for the final optimization. We further provide substantial experimental evidence that such a metric highly correlates with the real attack performance. We also consider more rigorous and realistic scenarios where the victims may hold high-resolution images and large-sized batches for training, and evaluate advanced defense strategies. Extensive experiments validate that GI-NAS can achieve state-of-the-art performance compared to existing gradient inversion methods. To the best of our knowledge, we are the first to introduce Neural Architecture Search (NAS) to gradient inversion attacks. Our main contributions are as follows:
  • We systematically analyze existing methods, emphasize the necessity of equipping image batch recovery with the optimal model structure, and propose GI-NAS to boost gradient inversion through neural architectural search.
  • We expand the model search space by considering different upsampling units and skip connection modes, and utilize a training-free search that regards the initial gradient matching loss as the search metric. We also provide extensive experimental evidence that such a metric highly correlates with the real attack performance.
  • Numerous experimental results demonstrate that GI-NAS outperforms state-of-the-art gradient inversion methods, even under extreme settings with high-resolution images, large-sized batches, and advanced defense strategies.
  • We provide deeper analysis on various aspects, including time costs, robustness to network parameters initialization, robustness to latent codes initialization, implications behind searched results, ablation study on each search type, and generalizability to more FL global models.
II Related Work
II-A Gradient Inversion Attacks and Defenses
Zhu et al. [12] first propose to restore the private samples via iterative optimization for pixel-level reconstruction, yet limited to low-resolution and single images. Geiping et al. [14] empirically decompose the gradient vector by its norm magnitude and updating direction, and succeed on high-resolution ImageNet [23] through the angle-based loss design. Furthermore, Yin et al. [15] extend the attacks to the batch-level inversion through the group consistency regularization and the improved batch label inference algorithm [13]. With strong handcrafted explicit priors (e.g., fidelity regularization, BN statistics), they accurately realize batch-level reconstruction with detailed semantic features allocated to all the individual images in a batch. Subsequent studies [16, 17, 18] leverage pre-trained GAN models as generative priors [24, 25, 26] to enhance the attacks. Current defenses focus on gradient perturbation to alleviate the impact of gradient inversion with degraded gradients [27, 28, 29]. Gaussian Noise [30] and Gradient Clipping [31] are common techniques in Differential Privacy (DP) [32, 33] that effectively constrain the attackers from learning through the released gradients. Gradient Sparsification [34, 35] prunes the gradients through a given threshold, and Soteria [36] edits the gradients from the aspect of learned representations.
II-B Neural Architecture Search (NAS)
By automatically searching the optimal model architecture, NAS algorithms prove to be significantly effective in multiple visual tasks such as image restoration [37, 38], semantic segmentation [39, 40, 41], and image classification [42, 43, 44, 45, 46]. In image classification tasks, Zoph et al. [43, 45] regularize the search space to a convolutional cell and construct a better architecture stacked by these cells using an RNN [47] controller. For semantic segmentation tasks, Liu et al. [40] search for the optimal network on the level of hierarchical network architecture and extend NAS to dense image prediction. In terms of image restoration tasks, Suganuma et al. [37] exploit a better Convolutional Autoencoders (CAE) with standard network components through the evolutionary search, while Chu et al. [38] discover a competitive lightweight model via both micro and macro architecture search.
Existing NAS methods adopt different strategies to exploit the search space, including evolutionary methods [48, 49, 50, 51, 52, 53, 54], Bayesian optimization [55, 56, 57], Reinforcement Learning (RL) [58, 59, 60, 43, 45], and gradient-based search [61, 62, 63, 64]. Different RL-based methods vary in the way to represent the agent’s policy and the optimization process. Zoph et al. [43] utilize RNN networks to sequentially encode the neural architecture and train them with the REINFORCE policy gradient algorithm [65]. Baker et al. [58] adopt Q-learning [66, 67] to train the policy network and realize competitive model design. Notably, recent works [68, 69, 70] propose training-free NAS to mitigate the issue of huge computational expenses. Instead of training from scratch, they evaluate the searched networks by some empirically designed metrics that reflect model effectiveness. In this paper, we adopt the initial gradient matching loss as the training-free search metric and provide substantial empirical evidence that such a search metric highly correlates with the real attack performance.
III Problem Formulation
Basics of Gradient Inversion. We consider the training process of a classification model $f_{\theta}$ parameterized by $\theta$ in FL scenarios. The real gradients $\mathbf{g}$ are calculated from a private batch (with real images $\mathbf{x}$ and real labels $\mathbf{y}$) at the client side. The universal goal of gradient inversion attacks is to search for some fake images $\mathbf{\hat{x}}\in\mathbb{R}^{B\times H\times W\times C}$ with labels $\mathbf{\hat{y}}\in\{0,1\}^{B\times L}$ so that ($\mathbf{\hat{x}}$, $\mathbf{\hat{y}}$) can be close to ($\mathbf{x}$, $\mathbf{y}$) as much as possible, where $B$, $H$, $W$, $C$, and $L$ are respectively batch size, image height, image width, number of channels, and number of classes. This can be realized by minimizing the gradient matching loss [12]:
$\mathbf{\hat{x}^{*}},\mathbf{\hat{y}^{*}}=\mathop{\arg\min}_{\mathbf{\hat{x}},\mathbf{\hat{y}}}\mathcal{D}(\nabla_{\theta}\mathcal{L}(f_{\theta}(\mathbf{\hat{x}}),\mathbf{\hat{y}}),\mathbf{g}),$ (1)
where $\mathcal{D(\cdot,\cdot)}$ is the distance metric (e.g., $l_{2}$-norm loss, cosine-similarity loss) for the gradient matching loss and $\mathcal{L}(\cdot,\cdot)$ is the loss function of the global model $f_{\theta}$.
Previous works [13, 15] in this field have revealed that the ground truth labels $\mathbf{y}$ can be directly inferred from the uploaded gradients $\mathbf{g}$. Therefore, the formulation in (1) can be simplified as:
$\mathbf{\hat{x}^{*}}=\mathop{\arg\min}_{\mathbf{\hat{x}}}\mathcal{D}(F(\mathbf{\hat{x}}),\mathbf{g}),$ (2)
where $F(\mathbf{\hat{x}})=\nabla_{\theta}\mathcal{L}(f_{\theta}(\mathbf{\hat{x}}),\mathbf{\hat{y}})$ calculates the gradients of $f_{\theta}$ provided with $\mathbf{\hat{x}}$.
The key challenge of (2) is that gradients only provide limited information of private data and there even exists a pair of different data having the same gradients [71]. To mitigate this issue, subsequent works incorporate various regularizations (e.g., total variation loss [14], group consistency loss [15]) as prior knowledge. Thus, the overall optimization becomes:
$\mathbf{\hat{x}^{*}}=\mathop{\arg\min}_{\mathbf{\hat{x}}}\mathcal{D}(F(\mathbf{\hat{x}}),\mathbf{g})+\lambda\mathcal{R}_{prior}(\mathbf{\hat{x}}),$ (3)
where $\mathcal{R}_{prior}(\cdot)$ is the introduced regularization that can establish some priors for the attacks, and $\lambda$ is the weight factor.
GAN-based Gradient Inversion. Nevertheless, the optimization of (3) is still limited in the pixel space. Given a well pre-pretrained GAN, an instinctive idea is to swift the optimization from the pixel space to the GAN latent space:
$\mathbf{z^{*}}=\mathop{\arg\min}_{\mathbf{z}}\mathcal{D}(F(G_{\omega}(\mathbf{z})),\mathbf{g})+\lambda\mathcal{R}_{prior}(G_{\omega}(\mathbf{z})),$ (4)
where $G_{\omega}$ and $\mathbf{z}\in\mathbb{R}^{B\times l}$ are respectively the generator and the latent vector of the pre-trained GAN. By reducing the optimization space from $\mathbb{R}^{B\times H\times W\times C}$ to $\mathbb{R}^{B\times l}$, (4) overcomes the uncertainty of directly optimizing the extensive pixels and exploits the abundant prior knowledge encoded in the pre-trained GAN. Based on this, recently emerged GAN-based attacks [16, 17] explore various search strategies within the pre-trained GAN to utilize its expression ability.
Gradient Inversion via Over-parameterized Networks. But as previously mentioned, incorporating such GAN priors is often impractical in realistic scenarios where the distribution of $\mathbf{x}$ is likely to be mismatched with the training data of the pre-trained GAN. Furthermore, it has already been discovered in [21] that explicitly introducing regularization in (3) or (4) may not necessarily result in convergence towards $\mathbf{x}$, as even ground truth images can not guarantee minimal loss when $\mathcal{R}_{prior}(\cdot)$ is added. To mitigate these issues, Zhang et al. [21] propose to leverage an over-parameterized network as implicit prior knowledge:
$\mathbf{\phi^{*}}=\mathop{\arg\min}_{\mathbf{\phi}}\mathcal{D}(F(G_{over}(\mathbf{z_{0}};\mathbf{\phi})),\mathbf{g}),$ (5)
where $G_{over}$ is the over-parameterized convolutional network with excessive parameters $\mathbf{\phi}$, and $\mathbf{z_{0}}$ is the randomly generated but fixed latent code. Note that the regularization term is omitted in (5). This is because the architecture of $G_{over}$ itself can serve as implicit regularization, for convolutional networks have been proven to possess implicit priors that prioritize clean images rather than noise as shown in [22]. Thus, the generated images that highly resemble the ground truth images can be obtained through $\mathbf{\hat{x}^{*}}=G_{over}(\mathbf{z_{0}};\mathbf{\phi^{*}})$. However, only a changeless over-parameterized network is employed for all the attack settings in [21]. As previously shown in Figure 1, although the network is over-parameterized, the attack performance exhibits significant differences when adopting different architectures. To fill this gap, we propose to further exploit such implicit architectural priors by searching the optimal over-parameterized network $G_{opt}$ for each batch. We will discuss how to realize this in Section IV.
IV Method
Figure 2: Overview of the proposed GI-NAS attack. We leverage a two-stage strategy for private batch recovery. In the first stage, we traverse the model search space and calculate the initial gradient matching loss (i.e., our training-free metric) of each model based on the fixed input $\mathbf{z_{0}}$. We regard the model that achieves the minimal initial loss as our best model, for its performance at the start can stand from numerous candidates. In the second stage, we adopt the architecture of the previously found best model and optimize its excessive parameters to reconstruct the private data.
Our proposed GI-NAS attack is carried out in two stages. In the first stage, we conduct our architecture search to decide the optimal model $G_{opt}$. Given that the attackers may only hold limited resources, we utilize the initial gradient matching loss as the training-free search metric to reduce the computational overhead. In the second stage, we iteratively optimize the parameters of the selected model $G_{opt}$ to recover the sensitive data. Figure 2 illustrates the overview of our method.
IV-A Training-free Optimal Model Search
IV-A1 Model Search Space Design
One crucial factor for NAS is that the potential search space is large and diverse enough to cover the optimal design. Therefore, we adopt U-Net [72], a typical convolutional neural architecture as the fundamental of our model search, since the skip connection patterns between its encoders and decoders can provide adequate alternatives of model structure. Besides, the configurations of the upsampling modules (e.g., kernel size, activation function) can also enable numerous possibilities when arranged and combined. Similar to previous NAS methods [73, 74], we enlarge the search space of our model from two aspects, namely Upsampling Modules and Skip Connection Patterns.
Search Space for Upsampling Modules. We decompose the upsampling operations into five key components: feature upsampling, feature transformation, activation function, kernel size, and dilation rate. Then, we allocate a series of possible options to each of these components. When deciding on feature upsampling, we choose from commonly used interpolation techniques, such as bilinear interpolation, bicubic interpolation, and nearest-neighbour interpolation. As for feature transformation, we choose from classical convolution techniques, such as 2D convolution, separable convolution, and depth-wise convolution. As regards activation function, we select from ReLU, LeakyReLU, PReLU, etc. Furthermore, we supply kernel size and dilation rate with more choices, such as $1\times 1$, $3\times 3$, or $5\times 5$ for kernel size and $1$, $3$, or $5$ for dilation rate. The combination of these flexible components can contribute to the diversity of upsampling modules.
Search Space for Skip Connection Patterns. We assume that there are $t$ levels of encoders and decoders in total, and denote them as $e_{1},e_{2},...,e_{t}$ and $d_{1},d_{2},...,d_{t}$. As shown in Figure 3, We consider different skip connection patterns between encoders and decoders. To represent each of these patterns, we define a skip connection matrix $\mathbf{A}\in\{0,1\}^{t\times t}$ that serves as a mask to determine whether there will be new residual connections [75] between pairs of encoders and decoders. More concretely, $\mathbf{A}_{ij}=1$ indicates that there exists a skip connection from $e_{i}$ to $d_{j}$ and $\mathbf{A}_{ij}=0$ means that there isn’t such a skip connection. As the shapes of feature maps across different network levels can vary significantly (e.g., $64\times 64$ for the output of $e_{i}$ and $256\times 256$ for the input of $d_{j}$), we introduce connection scale factors to tackle this inconsistency and decompose all the possible scale factors into a series of 2×2$\times$2 × upsampling operations or downsampling operations with shared weights. By allocating 00 or $1$ to each of the $t^{2}$ bits in A, we broaden the search space of skip connection patterns as there are $2^{t^{2}}$ possibilities altogether and we only need to sample a portion of them.
Figure 3: The design of search space for skip connection patterns. Different skip connection patterns are determined by the skip connection matrix $\mathbf{A}\in\{0,1\}^{t\times t}$. $\mathbf{A}_{ij}=1$ indicates that there exists a skip connection from $e_{i}$ to $d_{j}$ and $\mathbf{A}_{ij}=0$ means that there isn’t such a skip connection.
IV-A2 Optimal Model Selection
We build up our model search space $\mathcal{M}$ by combining the possibilities of the aforementioned upsampling modules and skip connection patterns. We assume that the size of our model search space is $n$ and the candidates inside it are denoted as =$\mathcal{M}=\{G_{1},G_{2},...,G_{n}\}$. We first sample the latent code $\mathbf{z_{0}}$ from the Gaussian distribution and freeze its values on all the models for fair comparison. We then traverse the model search space and calculate the initial gradient matching loss of each individual $G_{r}\,(1\leq r\leq n)$:
$\mathcal{L}_{grad}(G_{r})=\mathcal{D}(\mathcal{T}(F(G_{r}(\mathbf{z_{0}};\mathbf{\phi}_{r}))),\mathbf{g}),$ (6)
where $\mathcal{L}_{grad}(\cdot)$ is the gradient matching loss, $\mathcal{T}(\cdot)$ is the estimated gradient transformation [16], and $\mathbf{\phi}_{r}$ are the parameters of $G_{r}$. Here we introduce $\mathcal{T}(\cdot)$ to estimate the gradient transformation following the previous defense auditing work [16], since the victims may apply defense strategies to $\mathbf{g}$ (e.g., Gradient Clipping [30]) and only release disrupted forms of gradients. Empirically, the model that can perform the best at the start and stand out from numerous candidates is likely to have better implicit architectural priors with respect to the private batch. Therefore, we regard the model that achieves the minimal initial loss as our best model $G_{opt}$ and update our selection during the traversal. Since only the initial loss is calculated and no back-propagation is involved, this search process is training-free and hence computationally efficient.
Algorithm 1 Gradient Inversion via Neural Architecture Search (GI-NAS)
1:$n$: the size of the model search space; $\mathbf{g}$: the uploaded real gradients; $m$: the maximum iteration steps for the final optimization;
2:$\mathbf{\hat{x}^{*}}$: the generated images;
3:Build up the model search space: =$\mathcal{M}=\{G_{1},G_{2},...,G_{n}\}$ with the parameters $\Phi=\{\phi_{1},\phi_{2},...,\phi_{n}\}$
4:$\mathbf{z_{0}}\leftarrow\mathcal{N}(0,1)$, $loss_{min}\leftarrow+\infty$
5:for $i\leftarrow 1$ to $n$ do
6:     $loss_{i}\leftarrow\mathcal{D}(\mathcal{T}(F(G_{i}(\mathbf{z_{0}};\mathbf{\phi}_{i}))),\mathbf{g})$ // calculate the initial gradient matching loss
7:     if $loss_{i}<loss_{min}$ then
8:         $loss_{min}\leftarrow loss_{i}$, $G_{opt}\leftarrow G_{i}$, $\gamma_{1}\leftarrow\mathbf{\phi}_{i}$ // update the selection of $G_{opt}$
9:     end if
10:end for
11:for $k\leftarrow 1$ to $m$ do
12:     $\mathbf{\gamma}_{k+1}\leftarrow\mathbf{\gamma}_{k}-\eta\nabla_{\mathbf{\gamma}_{k}}\mathcal{D}(\mathcal{T}(F(G_{opt}(\mathbf{z_{0}};\mathbf{\gamma}_{k}))),\mathbf{g})$
13:end for
14:$\mathbf{\gamma^{*}}\leftarrow\mathbf{\gamma}_{m+1}$
15:return $\mathbf{\hat{x}^{*}}=G_{opt}(\mathbf{z_{0}};\mathbf{\gamma^{*}})$
IV-B Private Batch Recovery via the Optimal Model
After deciding on the optimal model $G_{opt}$ with the parameters $\gamma_{1}$, we iteratively conduct gradient decent optimizations on $G_{opt}$ to minimize the gradient matching loss:
$\mathbf{\gamma}_{k+1}=\mathbf{\gamma}_{k}-\eta\nabla_{\mathbf{\gamma}_{k}}\mathcal{L}_{grad}(G_{opt}),$ (7)
where $k$ is the current number of iterations $(1\leq k\leq m)$, $\gamma_{k+1}$ are the parameters of $G_{opt}$ after the $k$-th optimization, and $\eta$ is the learning rate. Once the above process converges and we obtain $\mathbf{\gamma^{*}}=\mathbf{\gamma}_{m+1}$ that satisfy the minimum loss, the private batch can be reconstructed by $\mathbf{\hat{x}^{*}}=G_{opt}(\mathbf{z_{0}};\mathbf{\gamma^{*}})$. In summary, the pseudocode of GI-NAS is illustrated in Algorithm 1.
V Experiments
V-A Experimental Setup
Evaluation Settings. We test our method on CIFAR-10 [76] and ImageNet [23] with the resolution of $32\times 32$ and $256\times 256$. Unlike many previous methods [17, 18] that scale down the ImageNet images to $64\times 64$, here we emphasize that we adopt the high-resolution version of ImageNet. Thus, our settings are more rigorous and realistic. Following previous gradient inversion works [12, 21], we adopt ResNet-$18$ [75] as our global model and utilize the same preprocessing procedures. We build up the search space and randomly generate the alternative models by arbitrarily changing the options of upsampling modules and skip connection patterns.
State-of-the-art Baselines for Comparison. We implement the following gradient inversion baselines: (1) IG (Inverting Gradients) [14]: pixel-level reconstruction with angle-based loss function; (2) GI (GradInversion) [15]: realizing batch-level restoration via multiple regularization priors; (3) GGL (Generative Gradient Leakage) [16]: employing strong GAN priors to produce high-fidelity images under severe defense strategies; (4) GIAS (Gradient Inversion in Alternative Space) [17]: searching the optimal latent code while optimizing in the generator parameter space; (5) GION (Gradient Inversion via Over-parameterized Networks) [21]: designing an over-parameterized convolutional network with excessive parameters and employing a fixed network architecture as implicit regularization. Note that when implementing GAN-based methods such as GGL and GIAS, we adopt BigGAN [77] pre-trained on ImageNet, which may result in mismatched priors when the target data is from CIFAR-10. However, such mismatch can be very common in realistic FL scenarios.
Quantitative Metrics. We utilize four metrics to measure the quality of reconstruction images: (1) PSNR $\uparrow$ (Peak Signal-to-Noise Ratio); (2) SSIM $\uparrow$ (Structural Similarity Index Measure); (3) FSIM $\uparrow$ (Feature Similarity Index Measure); (4) LPIPS $\downarrow$ (Learned Perceptual Image Patch Similarity) [78]. Note that $\downarrow$” indicates that the lower the metric, the better the attack performance while $\uparrow$” indicates that the higher the metric, the better the attack performance.
Implementation Details. The learning rate $\eta$ in (7) is set as $1\times 10^{-3}$. We utilize the signed gradient descent and adopt Adam optimizer [79] when updating the parameters of $G_{opt}$. We choose the negative cosine similarity function as the distance metric $\mathcal{D(\cdot,\cdot)}$ when calculating the gradient matching loss $\mathcal{L}_{grad}(\cdot)$ in (6). We conduct all the experiments on NVIDIA GeForce RTX 3090 GPUs for smaller batch sizes and on A100 GPUs for larger batch sizes.
V-B Choices of search size $n$ and network depth $t$
TABLE I: Attack results under different values of search size $n$. We randomly select images (strictly having no intersections with the images used in the remaining experiments of this paper) from ImageNet with the default batch size $B=4$ for testing.
Metric 1 10 50 100 500 1000 2000 3000 4000 5000 6000 7000 8000 9000 PSNR $\uparrow$ 21.5034 21.6179 21.2776 22.0846 22.1451 22.5949 22.7380 23.3836 23.5626 24.1603 23.1646 23.3276 23.2485 22.9561 SSIM $\uparrow$ 0.6314 0.5960 0.6222 0.6069 0.6099 0.6335 0.5960 0.6314 0.6424 0.6566 0.6518 0.6395 0.6324 0.6306 FSIM $\uparrow$ 0.8281 0.7938 0.8019 0.8082 0.8125 0.8269 0.8028 0.8243 0.8337 0.8439 0.8334 0.8347 0.8319 0.8228 LPIPS $\downarrow$ 0.4391 0.5238 0.5212 0.4808 0.4839 0.4540 0.4499 0.4369 0.4317 0.4103 0.4289 0.4281 0.4282 0.4913
TABLE II: Attack results under different values of network depth $t$. We randomly select images (strictly having no intersections with the images used in the remaining experiments of this paper) from ImageNet with the default batch size $B=4$ for testing.
Metric 3 4 5 6 7 PSNR $\uparrow$ 21.3108 22.4132 24.1603 23.1865 22.9935 SSIM $\uparrow$ 0.6114 0.6314 0.6566 0.6441 0.6293 FSIM $\uparrow$ 0.8249 0.8314 0.8439 0.8317 0.8287 LPIPS $\downarrow$ 0.4478 0.4263 0.4103 0.4263 0.4301
To obtain the best reconstruction performance, we need to carefully decide on the values of search size $n$ and network depth $t$. Note that both these two hyper-parameters reflect the degree of NAS, and there is a trade-off between under-fitting and over-fitting when choosing their values. Thus, we randomly select images (strictly having no intersections with the images used in the remaining experiments of this paper) from ImageNet and test the attack results under different values of $n$ and $t$ with the default batch size $B=4$. From Table I and Table II, we find that when $n$ and $t$ are small, increasing any one of them indeed improves the attacks, as more optimal alternatives are provided or the models are more complicated to tackle the under-fitting. However, when $n$ and $t$ are large, further increasing any one of them cannot improve the attacks and will conversely lead to performance degradation due to the over-fitting. This is because candidates that focus too much on the minimal initial loss while neglecting the fundamental patterns of data recovery may be included in the search space and eventually selected. Thus, we adopt $n=5000$ and $t=5$ in the remaining experiments of this paper, as they strike a better balance and acquire the best results in Table I and Table II.
TABLE III: Quantitative comparison of GI-NAS to state-of-the-art gradient inversion methods on CIFAR-10 ($32\times 32$) and ImageNet ($256\times 256$) with the default batch size $B=4$.
Dataset Metric IG GI GGL GIAS GION GI-NAS CIFAR-10 PSNR $\uparrow$ 16.3188 15.4613 12.4938 17.3687 30.8652 35.9883 SSIM $\uparrow$ 0.5710 0.5127 0.3256 0.6239 0.9918 0.9983 FSIM $\uparrow$ 0.7564 0.7311 0.6029 0.7800 0.9960 0.9991 LPIPS $\downarrow$ 0.4410 0.4878 0.5992 0.4056 0.0035 0.0009 ImageNet PSNR $\uparrow$ 7.9419 8.5070 11.6255 10.0602 21.9942 23.2578 SSIM $\uparrow$ 0.0815 0.1157 0.2586 0.2408 0.6188 0.6848 FSIM $\uparrow$ 0.5269 0.5299 0.5924 0.5719 0.8198 0.8513 LPIPS $\downarrow$ 0.7194 0.7168 0.6152 0.6563 0.4605 0.3952
V-C Comparison with State-of-the-art Methods
Firstly, we compare GI-NAS with state-of-the-art gradient inversion methods on CIFAR-10. From Table III, we conclude that GI-NAS achieves the best results with significant performance improvement. For instance, we realize a $5.12$ dB PSNR increase than GION, and our LPIPS value is $74.3\%$ smaller than that of GION. These prove that in contrast to GION that optimizes on a fixed network, our NAS strategy indeed comes into effect and better leverages the implicit architectural priors.
TABLE IV: Quantitative comparison of GI-NAS to state-of-the-art gradient inversion methods on CIFAR-10 ($32\times 32$) and ImageNet ($256\times 256$) with larger batch sizes when $B>4$.
Dataset Metric Batch Size IG GI GGL GIAS GION GI-NAS Batch Size IG GI GGL GIAS GION GI-NAS CIFAR-10 PSNR $\uparrow$ 16 10.2647 11.2593 9.9964 11.1952 24.8746 30.5315 32 9.5802 10.5863 9.4104 9.8669 28.6832 33.0586 SSIM $\uparrow$ 0.2185 0.2431 0.1658 0.2638 0.9816 0.9948 0.1595 0.2020 0.1481 0.1724 0.9892 0.9974 FSIM $\uparrow$ 0.5254 0.5350 0.4988 0.5569 0.9869 0.9965 0.5186 0.5325 0.4995 0.5180 0.9882 0.9984 LPIPS $\downarrow$ 0.6254 0.6225 0.6655 0.5804 0.0177 0.0035 0.6598 0.6626 0.6922 0.6362 0.0218 0.0017 PSNR $\uparrow$ 48 9.5446 10.2000 9.1200 9.6376 29.5312 38.7054 96 9.2363 9.9500 8.8575 9.2982 28.9723 31.8026 SSIM $\uparrow$ 0.1574 0.1697 0.1552 0.1652 0.9932 0.9983 0.1402 0.1643 0.1516 0.2180 0.9878 0.9941 FSIM $\uparrow$ 0.5112 0.5167 0.4869 0.5126 0.9949 0.9987 0.4961 0.5238 0.4801 0.5365 0.9865 0.9941 LPIPS $\downarrow$ 0.6660 0.6571 0.7213 0.6554 0.0072 0.0020 0.6865 0.6593 0.7248 0.7297 0.0227 0.0071 ImageNet PSNR $\uparrow$ 8 7.6011 8.3419 11.5454 9.6216 20.4841 20.8451 16 7.7991 8.0233 11.4766 9.2563 20.3078 21.4267 SSIM $\uparrow$ 0.0762 0.1113 0.2571 0.2218 0.5681 0.6076 0.0704 0.0952 0.2561 0.2189 0.5507 0.6141 FSIM $\uparrow$ 0.5103 0.5365 0.5942 0.5765 0.7908 0.8105 0.5077 0.5165 0.5872 0.5832 0.7840 0.8157 LPIPS $\downarrow$ 0.7374 0.7251 0.6152 0.6621 0.5162 0.4613 0.7424 0.7362 0.6165 0.6656 0.5313 0.4613 PSNR $\uparrow$ 24 7.1352 7.9079 10.9910 9.6601 20.6845 21.7933 32 7.1085 7.8752 10.8987 10.0563 16.1275 20.0011 SSIM $\uparrow$ 0.0549 0.0892 0.2583 0.2270 0.5868 0.6236 0.0554 0.0763 0.2558 0.2361 0.3415 0.5485 FSIM $\uparrow$ 0.4794 0.4943 0.5811 0.5618 0.7978 0.8163 0.4940 0.4951 0.5889 0.5762 0.6823 0.7819 LPIPS $\downarrow$ 0.7540 0.7475 0.6261 0.6667 0.5219 0.4651 0.7493 0.7427 0.6296 0.6674 0.6291 0.5265
We also discover that the GAN-based method GGL underperforms the previous GAN-free methods (i.e., IG and GI). This is because GGL utilizes BigGAN [77] pre-trained on ImageNet for generative priors in our settings, which has an inherent distribution bias with the target CIFAR-10 data domain. Besides, GGL only optimizes the latent vectors and cannot dynamically handle the mismatch between the training data of GAN and the target data. In contrast, GIAS optimizes both the latent vectors and the GAN parameters, which can reduce the distribution divergence and alleviate such mismatch to some extent. Therefore, although also GAN-based, GIAS exhibits much better performance than GGL on CIFAR-10.
High-Resolution Images Recovery. We then consider a more challenging situation and compare various methods on ImageNet with the resolution of $256\times 256$. In Table III, most methods encounter significant performance decline when attacking high-resolution images. The amplification of image pixels greatly increases the complexity of reconstruction tasks and thus obstructs the optimization process for optimal images. But GI-NAS still achieves the best attack results, with a PSNR increase of $1.26$ dB than GION.
Original
IG
GI
GGL
GIAS
GION
GI-NAS
Figure 4: Qualitative comparison of GI-NAS to state-of-the-art gradient inversion methods on ImageNet ($256\times 256$) with the larger batch size $B=32$.
Extension to Larger Batch Sizes. As shown in Table IV, we extend GI-NAS to larger batch sizes both on CIFAR-10 and ImageNet, which is more in line with the actual training process of FL systems. We observe that the performance of most methods degrades as batch sizes increase because it would be harder for the attackers to distinguish each individuals in a batch, while GI-NAS is insusceptible to larger batch sizes and continuously generates high-quality images even at $B=96$ on CIFAR-10 and $B=32$ on ImageNet. Besides, GI-NAS is able to acquire consistent and significant performance gains on the basis of GION at all the given batch sizes. Figure 4 shows the qualitative comparison on ImageNet with the larger batch size $B=32$. We notice that GI-NAS exactly realizes pixel-level recovery, while all the other compared methods struggle to perform the attacks and only obtain images with huge visual differences from the original ones. These results further provide evidence for the necessity and effectiveness of our batch-level optimal architecture search.
Attacks under Defense Strategies. Next, we evaluate how these attacks perform when defenses are applied both on CIFAR-10 and ImageNet. Following previous works [16, 12], we consider four strict defense strategies: (1) Gaussian Noise [30] with a standard deviation of $0.1$; (2) Gradient Sparsification [34] with a pruning rate of $90\%$; (3) Gradient Clipping [30] with a clipping bound of $4$; (4) Representative Perturbation (Soteria) [36] with a pruning rate of $80\%$. For fair comparison, we apply the estimated gradient transformation $\mathcal{T}(\cdot)$ described in (6) to all the attack methods. From Table V, we discover that although the gradients have been disrupted by the imposed defense strategies, GI-NAS still realizes the best reconstruction effects in almost all the tested cases. The only exception is that GGL outperforms GI-NAS in terms of SSIM and LPIPS when the defense strategy is Gaussian Noise [30]. This is because the gaussian noise with a standard deviation of $0.1$ can severely corrupt the gradients and the information carried inside the gradients is no longer enough for private batch recovery. However, GGL only optimizes the latent vectors and can still generate natural images that possess some semantic features by the pre-trained GAN. Thus, GGL can still obtain not bad performance even though the generated images are quite dissimilar to the original ones.
TABLE V: Quantitative comparison of GI-NAS to state-of-the-art gradient inversion methods on CIFAR-10 ($32\times 32$) and ImageNet ($256\times 256$) under various defense strategies with the default batch size $B=4$.
Dataset Metric Defense IG GI GGL GIAS GION GI-NAS Defense IG GI GGL GIAS GION GI-NAS CIFAR-10 PSNR $\uparrow$ Gaussian Noise 8.8332 8.2797 10.3947 9.8146 9.7658 10.5192 Gradient Sparsification 9.6183 10.9486 10.4323 12.1328 30.3518 39.8739 SSIM $\uparrow$ 0.1177 0.0976 0.1856 0.1301 0.0280 0.0375 0.1592 0.2168 0.1862 0.3227 0.9920 0.9987 FSIM $\uparrow$ 0.4844 0.4806 0.5123 0.4843 0.5945 0.6300 0.5151 0.5329 0.5223 0.6015 0.9951 0.9987 LPIPS $\downarrow$ 0.7258 0.7077 0.6255 0.7129 0.6752 0.6740 0.6895 0.6855 0.6298 0.5582 0.0047 0.0014 PSNR $\uparrow$ Gradient Clipping 16.6203 11.5888 10.1373 17.9859 32.4744 36.2490 Soteria 9.4103 10.2939 10.4217 15.5825 31.1003 34.1804 SSIM $\uparrow$ 0.5952 0.2851 0.1760 0.6703 0.9933 0.9985 0.1404 0.1707 0.1889 0.5405 0.9900 0.9973 FSIM $\uparrow$ 0.7464 0.4976 0.5070 0.7751 0.9966 0.9992 0.5086 0.5086 0.5144 0.7133 0.9952 0.9986 LPIPS $\downarrow$ 0.3811 0.6413 0.6205 0.3054 0.0029 0.0008 0.7008 0.6940 0.6335 0.3748 0.0038 0.0014 ImageNet PSNR $\uparrow$ Gaussian Noise 7.7020 7.4553 9.7381 6.8520 9.2449 11.2567 Gradient Sparsification 8.6540 8.6519 9.8019 9.7894 13.7358 15.4331 SSIM $\uparrow$ 0.0311 0.0197 0.2319 0.1045 0.0277 0.0416 0.0669 0.0627 0.2315 0.2515 0.2387 0.3154 FSIM $\uparrow$ 0.4518 0.3804 0.5659 0.4562 0.5714 0.5846 0.5073 0.4696 0.5624 0.5664 0.6365 0.6786 LPIPS $\downarrow$ 0.7775 0.8220 0.6639 0.7282 0.8033 0.7805 0.7480 0.7681 0.6589 0.6604 0.6909 0.6535 PSNR $\uparrow$ Gradient Clipping 8.0690 7.2342 11.3215 9.7556 22.2016 23.2074 Soteria 6.5675 6.6693 11.5690 9.7588 22.9617 23.7917 SSIM $\uparrow$ 0.0881 0.0659 0.2531 0.2547 0.6266 0.7023 0.0323 0.0382 0.2623 0.2425 0.6698 0.7143 FSIM $\uparrow$ 0.5362 0.5434 0.5900 0.5794 0.8229 0.8635 0.4929 0.4454 0.5964 0.5697 0.8430 0.8658 LPIPS $\downarrow$ 0.7227 0.7230 0.6117 0.6485 0.4603 0.3743 0.7506 0.7703 0.6031 0.6631 0.4253 0.3655
TABLE VI: Correlation indicators between various training-free search metrics and the actual PSNR performance on CIFAR-10 ($32\times 32$) and ImageNet ($256\times 256$) with the default batch size $B=4$. Note that a correlation indicator closer to 1 (or -1) indicates a stronger positive (or negative) correlation.
Training-Free Search Metric CIFAR-10 ImageNet Kendall’s $\tau$ Pearson Coefficient Spearman Coefficient Kendall’s $\tau$ Pearson Coefficient Spearman Coefficient Gaussian Noise $\mathcal{N}\sim(0,1)$ 0.0531 0.0948 0.0720 0.1187 0.1672 0.1773 Uniform Noise $\mathcal{U}\sim(0,1)$ 0.0885 0.0921 0.1291 0.0486 0.0822 0.0798 Initial Gradient Matching Loss -0.4491 -0.6128 -0.6345 -0.4906 -0.6633 -0.7120
V-D Further Analysis
(a) CIFAR-10
(b) ImageNet
Figure 5: Correlation between the initial gradient matching loss and the actual PSNR performance on CIFAR-10 ($32\times 32$) and ImageNet ($256\times 256$) with the default batch size $B=4$.
Effectiveness of Our Training-Free Search Metric. We provide extensive empirical evidence that our training-free search metric (i.e., the initial gradient matching loss) highly negatively correlates with the real attack performance. Specifically, we attack the same private batch using models with different architectures and report the PSNR performance as well as the training-free search metric of each model both on CIFAR-10 and ImageNet. Following previous training-free NAS methods [80, 81, 82], we calculate the correlation indicators (i.e., Kendall’s $\tau$ [83], Pearson Coefficient [84], Spearman Coefficient [85]) to quantitatively measure the relevance. To be more convincing, we introduce Gaussian Noise $\mathcal{N}\sim(0,1)$ and Uniform Noise $\mathcal{U}\sim(0,1)$ for comparison. From Figure 5 and Table VI, we conclude that a smaller initial gradient matching loss is more likely to result in better PSNR performance, while the random noises are far from similar effects. Thus, utilizing the initial gradient matching loss as the training-free search metric is reasonable and meaningful.
TABLE VII: Time costs (minute) of different methods averaged over each batch on CIFAR-10 ($32\times 32$) and ImageNet ($256\times 256$) with the default batch size $B=4$.
Dataset Metric IG GI GGL GIAS GION GI-NAS CIFAR-10 NAS Time - - - - - 6.6 Total Time 46.1 32.9 131.4 214.8 46.9 57.3 PSNR 16.3188 15.4613 12.4938 17.3687 30.8652 35.9883 ImageNet NAS Time - - - - - 10.9 Total Time 291.1 273.7 244.6 343.2 58.1 84.4 PSNR 7.9419 8.5070 11.6255 10.0602 21.9942 23.2578
Computational Efficiency. We next analyze the computational costs. From Table VII, we find that NAS time only takes around $12\%$ of Total Time, yet GI-NAS still achieves significant PSNR gains over GION and reaches state-of-the-art. The additional time costs introduced by our training-free NAS ($6.6$ or $10.9$ minutes per batch) are very minor, especially given that many methods (e.g., GGL, GIAS) consume over $100$ or even $200$ minutes per batch. Therefore, such minor extra time costs in exchange for significant quantitative improvements are essential and worthwhile. Our method achieves a good trade-off between attack performance and computational efficiency.
TABLE VIII: Results when attacking the same batch by different networks using different parameter initialization seeds on CIFAR-10 ($32\times 32$) and ImageNet ($256\times 256$) with the default batch size $B=4$. Note that $\sigma_{loss}$ and $\sigma_{psnr}$ are respectively the standard deviations of the initial gradient matching losses and PSNR results on each network.
Dataset Model Random Seed Initial Gradient Matching Loss (×$\times 10^{-3}$) $\sigma_{loss}$ (×$\times 10^{-4}$) PSNR $\sigma_{psnr}$ CIFAR-10 $G_{c1}$ $s_{c1}$ 8.5647 0.0012 28.7283 0.1211 $s_{c2}$ 8.5648 28.8082 $s_{c3}$ 8.5645 28.5703 $G_{c2}$ $s_{c1}$ 8.4321 0.0006 33.5848 0.2581 $s_{c2}$ 8.4323 33.4631 $s_{c3}$ 8.4322 33.9584 $G_{c3}$ $s_{c1}$ 7.9501 0.0015 49.4290 0.2936 $s_{c2}$ 7.9499 49.1171 $s_{c3}$ 7.9502 49.7039 ImageNet $G_{i1}$ $s_{i1}$ 21.0167 5.6866 22.9013 0.0329 $s_{i2}$ 22.1401 22.8912 $s_{i3}$ 21.4250 22.8398 $G_{i2}$ $s_{i1}$ 24.4600 2.2089 21.1475 0.3040 $s_{i2}$ 24.0722 21.6449 $s_{i3}$ 24.0828 21.6989 $G_{i3}$ $s_{i1}$ 18.5319 6.0290 23.5278 0.1884 $s_{i2}$ 18.3819 23.2974 $s_{i3}$ 17.4208 23.6709
Robustness to Network Parameters Initialization. Following the general paradigm of previous training-free NAS methods [80, 81, 82] in other areas, the parameters of all the NAS candidates are randomly initialized and there is no special need to tune the parameters before calculating the initial gradient matching losses. Moreover, GI-NAS records both the architectures as well as the initial parameters when finding the optimal models. Therefore, for the model we decide on, its initial gradient matching loss in the NAS time (i.e., Stage 1 as shown in Figure 2) is the same as its initial gradient matching loss in the optimization time (i.e., Stage 2 as shown in Figure 2), and the influences of parameters initialization can be further reduced. To further demonstrate the robustness, we randomly select models with different architectures and apply different parameter initialization seeds to them to attack the same batch. From Table VIII, we find that for the same architecture, both the initial gradient matching losses and the final PSNR results stay within a steady range when changing the parameter initialization seeds. However, when the architectures change, both the initial gradient matching losses and the final PSNR results will vary significantly. The impacts of different architectures are much greater than the impacts of different parameter initialization seeds. As a result, we conclude that GI-NAS is stable and insensitive to network parameters initialization. We also observe that the models with smaller initial gradient matching losses often have better final PSNR results in Table VIII, which is consistent with our aforementioned findings in Figure 5 and Table VI. This again validates the rationality and usefulness of adopting the initial gradient matching loss as the training-free search metric.
Robustness to Latent Codes Initialization. The latent code $\mathbf{z_{0}}$ is randomly sampled from the Gaussian distribution $\mathcal{N}(0,1)$ and then kept frozen. Thus, we explore the influences of different latent codes by changing their initialization seeds. From Table IX, we discover that although different latent code initialization seeds on the same architecture may result in different initial gradient matching losses, the differences are still very minor (less than $10^{-6}$ for CIFAR-10 and less than $10^{-3}$ for ImageNet), and the impacts on the final PSNR results are limited (less than $0.5$ dB for CIFAR-10 and less than $0.4$ dB for ImageNet). However, when using different architectures, both these two values will change dramatically, and the differences are much greater than those brought by changing the latent code initialization seeds. Hence, GI-NAS is also robust to latent codes initialization.
TABLE IX: Results when attacking the same batch by different networks using different latent code initialization seeds on CIFAR-10 ($32\times 32$) and ImageNet ($256\times 256$) with the default batch size $B=4$. Note that $\sigma_{loss}$ and $\sigma_{psnr}$ are respectively the standard deviations of the initial gradient matching losses and PSNR results on each network.
Dataset Model Random Seed Initial Gradient Matching Loss (×$\times 10^{-3}$) $\sigma_{loss}$ (×$\times 10^{-4}$) PSNR $\sigma_{psnr}$ CIFAR-10 $G_{c4}$ $s_{c4}$ 14.5338 0.0018 26.4506 0.1925 $s_{c5}$ 14.5335 26.2082 $s_{c6}$ 14.5336 26.0703 $G_{c5}$ $s_{c4}$ 13.4751 0.0018 30.1240 0.2675 $s_{c5}$ 13.4753 30.2045 $s_{c6}$ 13.4755 30.6222 $G_{c6}$ $s_{c4}$ 13.2704 0.0015 31.3139 0.0566 $s_{c5}$ 13.2701 31.4256 $s_{c6}$ 13.2702 31.3540 ImageNet $G_{i4}$ $s_{i4}$ 23.1931 4.6280 21.8376 0.1164 $s_{i5}$ 23.6335 22.0556 $s_{i6}$ 22.7083 21.8760 $G_{i5}$ $s_{i4}$ 19.5264 1.1102 23.6033 0.1889 $s_{i5}$ 19.5677 23.2439 $s_{i6}$ 19.7360 23.3226 $G_{i6}$ $s_{i4}$ 25.9479 3.6054 20.4730 0.1655 $s_{i5}$ 25.3497 20.8007 $s_{i6}$ 25.9975 20.6776
TABLE X: Averaged architectural statistics of the searched optimal models by our NAS strategy under various settings tested in Section V-C. The default settings are highlighted in gray color. $\Delta_{connection}$ and $\Delta_{upsampling}$ are respectively the gains of the number of skip connections and the kernel size of upsampling modules when compared with the default settings.
Dataset Batch Size Defense Number of Skip Connections Kernel Size of Upsampling Modules $\Delta_{connection}$ $\Delta_{upsampling}$ Mostly Used Upsampling Operations CIFAR-10 4 No Defense 11.4 4.8 0.0 0.0 bilinear 16 No Defense 13.5 4.0 2.1 -0.8 bilinear 32 No Defense 14.0 5.0 2.6 0.3 bicubic 48 No Defense 13.0 5.0 1.6 0.3 bicubic 96 No Defense 12.0 7.0 0.6 2.3 bicubic 4 Gaussian Noise 12.3 5.5 0.9 0.8 bilinear 4 Gradient Sparsification 11.8 4.5 0.4 -0.3 pixel shuffle 4 Gaussian Clipping 12.5 5.0 1.1 0.3 bilinear 4 Soteria 11.5 5.0 0.1 0.3 bicubic ImageNet 4 No Defense 11.9 3.8 0.0 0.0 bilinear 8 No Defense 12.8 5.5 0.9 1.8 bilinear 16 No Defense 14.0 6.0 2.1 2.3 bilinear 24 No Defense 16.0 5.0 4.1 1.3 bilinear 32 No Defense 12.0 7.0 0.1 3.3 bicubic 4 Gaussian Noise 12.3 5.0 0.4 1.3 pixel shuffle 4 Gradient Sparsification 12.4 4.3 0.5 0.5 bilinear 4 Gaussian Clipping 14.1 3.3 2.3 -0.5 bilinear 4 Soteria 12.6 4.3 0.8 0.5 bilinear
Preferences of NAS Outcomes. To figure out the implications behind NAS preferences, we trace back the searched optimal models from Section V-C and analyze their averaged architectural statistics in Table X. Note that the bicubic operation takes more neighbourhood pixels than the bilinear operation ($16>4$) and is thus considered to be more complex. Besides, the pixel shuffle operation rearranges lots of pixels and is also considered to be more complicated than the bilinear operation. From Table X, we observe that the optimal models are largely dependent on recovery difficulties. Generally, batches with larger sizes or under defenses are more difficult to recover. Compared to the default settings (highlighted in gray color) that are with the batch size $B=4$ and under no defenses, the more difficult settings would prefer models with higher complexity (e.g., more skip connections, larger kernel sizes, more complex operations). The complexity gains $\Delta_{connection}$ and $\Delta_{upsampling}$ are often greater than 00 when the settings are $B>4$ or under defenses. Easier batches prefer simpler, lighter architectures and harder batches prefer more complex, heavier architectures. This might be used to further improve the search efficiency. By considering the attack difficulties ahead of time, we can have a better estimation of the complexity of optimal models. Thus, we only need to focus on models with a certain type of complexity and skip other unrelated models even when there is no enough time for NAS.
Ablation Study. In Table XI, We report the PSNR results of GION and different variants of GI-NAS. GI-NAS optimizes on a fixed over-parameterized network with the search size $n=1$, which means that it is essentially the same as GION. GI-NAS only searches the skip connection patterns while GI-NAS only searches the upsampling modules. We observe that the performance of GI-NAS is very close to that of GION, as both of them adopt a changeless architecture. GI-NAS and GI-NAS indeed improve the attacks on the basis of GION or GI-NAS, which validates the contributions of individual search types. GI-NAS combines the above two search types and thus performs the best among all the variants.
TABLE XI: Ablation Study. We report the PSNR results of GION and different variants of GI-NAS on CIFAR-10 ($32\times 32$) and ImageNet ($256\times 256$) with the default batch size $B=4$.
Method Connection Search Upsampling Search CIFAR-10 ImageNet GION 30.8652 21.9942 GI-NAS 30.6514 22.0126 GI-NAS 35.4336 22.4597 GI-NAS 35.3542 22.3992 GI-NAS 35.9883 23.2578
Generalizability to More FL Global Models. We show the performance of different methods when attacking more FL global models (e.g., LeNet-Zhu [12], ConvNet-32 [14]) in Figure 6. We note that GI-NAS realizes the best reconstruction results in all the tested cases. These results further demonstrate the reliability of GI-NAS. We also find that the attack results on different FL global models are distinctively dissimilar even when applying the same gradient inversion method. This implies that future study may have a deeper look into the correlation between the gradient inversion robustness and the architecture of FL global model, and thus design more securing defense strategies from the perspective of adaptive network architecture choice.
(a) CIFAR-10
(b) ImageNet
Figure 6: PSNR results of GI-NAS and state-of-the-art gradient inversion methods when attacking more FL global models on CIFAR-10 ($32\times 32$) and ImageNet ($256\times 256$) with the default batch size $B=4$.
VI Conclusion
In this paper, we propose GI-NAS, a novel gradient inversion method that makes deeper use of the implicit architectural priors for gradient inversion. We first systematically analyze existing gradient inversion methods and emphasize the necessity of adaptive architecture search. We then build up our model search space by designing different upsampling modules and skip connection patterns. To reduce the computational overhead, we leverage the initial gradient matching loss as the training-free search metric to select the optimal model architecture and provide extensive experimental evidence that such a metric highly correlates with the real attack performance. Extensive experiments prove that GI-NAS can achieve state-of-the-art performance compared to existing methods, even under more practical FL scenarios with high-resolution images, large-sized batches, and advanced defense strategies. We also provide deeper analysis on various aspects, including time costs, robustness to network parameters initialization, robustness to latent codes initialization, implications behind searched results, ablation study on each search type, and generalizability to more FL global models. We hope that the remarkable attack improvements of GI-NAS over existing gradient inversion methods may help raise the public awareness of such privacy threats, as the sensitive data would be more likely to be revealed or even abused. Moreover, we hope that the idea of this paper may shed new light on the gradient inversion community and facilitate the research in this field.
Limitations and Future Directions. Although we have empirically demonstrated that the initial gradient matching loss is highly negatively correlated to the final attack performance, it is still of great significance to further prove the effectiveness of this search metric with rigorous theoretical analysis. In the future work, we hope to provide more insightful theoretical results with regard to this search metric from various perspectives, such as the frequency spectrum [86, 74] or the implicit neural architectural priors [22, 87].